# GraphUniverse: Enabling Systematic Evaluation of Inductive Generalization

### v1

@@ Abstract
A fundamental challenge in graph learning is understanding how models generalize to new, unseen graphs. While synthetic benchmarks offer controlled settings for analysis, existing approaches are confined to single-graph, transductive settings where models train and test on the same graph structure. Addressing this gap, we introduce GraphUniverse, a framework for generating entire families of graphs to enable the first systematic evaluation of inductive generalization at scale. Our core innovation is the generation of graphs with persistent semantic communities, ensuring conceptual consistency while allowing fine-grained control over structural properties like homophily and degree distributions. This enables crucial but underexplored robustness tests, such as performance under controlled distribution shifts. Benchmarking a wide range of architectures—from GNNs to graph transformers and topological architectures—reveals that strong transductive performance is a poor predictor of inductive generalization. Furthermore, we find that robustness to distribution shift is highly sensitive not only to model architecture choice but also to the initial graph regime (e.g., high vs. low homophily). Beyond benchmarking, GraphUniverse’s flexibility and scalability can facilitate the development of robust and truly generalizable architectures—including next-generation graph foundation models. An interactive demo is available at https://graphuniverse.streamlit.app/.
@@ 1 Introduction
Graph learning has emerged as a powerful paradigm for learning from relational data across diverse domains, from drug discovery (Wong et al., 2024) and fraud detection (Cheng et al., 2025) to knowledge graphs (Galkin et al., 2023). Graph Neural Networks (GNNs) (Scarselli et al., 2008), with its countless variants (Gilmer et al., 2017; Kipf, 2016; Hamilton et al., 2017; Xu et al., 2018; Veličković et al., 2017), have demonstrated remarkable success in extending deep learning frameworks to graph-structured data, achieving competitive performance on tasks ranging from node classification to graph-level prediction. However, unlike the transformative leap from task-specific models to general-purpose architectures observed in natural language processing and computer vision, graph learning remains largely limited to specialized, task-specific models with limited evidence of robust generalization and scaling capabilities.
@@
    We develop a hierarchical generative model extending Degree Corrected-Stochastic Block Models (DC-SBMs) (Karrer & Newman, 2011) to an inductive setting with multiple graphs that maintain semantic consistency—i.e. node identities or community structures persist across different graph instances—while enabling controlled variation in their structural properties.
@@
  3. 3.
    We conduct systematic benchmarking comparing inductive and transductive evaluation across classical and contemporary graph architectures, revealing differences in model rankings between paradigms. Additionally, we evaluate model robustness under controlled property shifts, an analysis only possible with our inductive framework, finding that robustness strongly depends on both architecture choice and initial graph properties. These findings challenge conventional assumptions about graph model performance and demonstrate the critical importance of evaluation paradigm choice in assessing true model capabilities.
  4. 4.
    Upon acceptance, we will publicly release GraphUniverse as a Python package via PyPi for programmatic use, as well as publish its full implementation into TopoBench (Telyatnikov et al., 2024) to easily reproduce and/or expand our experimental results.
We envision GraphUniverse as a versatile tool for diverse research applications, from targeted generalization benchmarks to large-scale data generation for pre-training graph foundation models. While our experiments demonstrate its immediate utility, they represent only a fraction of what is possible with controllable graph family generation. Thus, we release GraphUniverse to the community as a flexible framework, inviting extensions and adaptations to explore new frontiers in graph learning.
@@ 2 Related Work
Recognizing the limitations of static datasets, a growing consensus advocates for high-fidelity synthetic data generation as a path toward more principled and scalable evaluation (Bechler-Speicher et al., 2025; Wang et al., 2025). The most prominent effort in this direction is GraphWorld (Palowitch et al., 2022), which uses synthetic generation to study model performance across a space of graph properties. While this allows for controlled analysis, GraphWorld’s methodology is confined to generating independent, single graphs. This restricts evaluation to the transductive setting, where models are tested on the same graph structure seen during training, and fundamentally prevents the study of a model’s ability to generalize to entirely new and unseen graphs.
GraphUniverse directly addresses the limitations of both prior paradigms. Unlike static benchmarks like OGB and GOOD, it provides a generative framework capable of producing unlimited data with fine-grained, tunable control over structural properties. And critically, unlike GraphWorld, it generates entire families of graphs with shared semantic meaning, enabling the first systematic, large-scale study of inductive generalization. This allows for the precise measurement of model robustness under controlled distribution shifts and provides the scalable data necessary to develop the next generation of truly generalizable graph learning models.
@@ 3 Background
A limitation of SBM is that it enforces homogeneous expected degrees within a community, since edge probabilities depend only on block membership. To address this, the original DC-SBM (Karrer & Newman, 2011) implementation introduces node-specific degree factors $\theta_{i}>0$ and a nonnegative block matrix $\Lambda\in\mathbb{R}_{\geq 0}^{k\times k}$. Moreover, they shift focus from simple graphs to multigraphs, i.e. a graph with multi-edges and self-loops. In its original Poisson multigraph form, edges are counts with rates:
@@
so $\Lambda$ controls the total number of edges in a block, while $\theta$ redistributes degrees within blocks. This Poisson formulation is standard in inference-focused works (Karrer & Newman, 2011; Abbe, 2017).
@@
Since the original DC-SBM naturally generate Poisson multigraphs, a common approach to get simple graphs from them is to collapse multi-edges into single undirected edges and remove self-edges after generation (as done in GraphWorld (Palowitch et al., 2022)). However, this leads to a systematic but unpredictable mismatch between the input parameters and the properties of the output graph. Therefore, we choose to work directly with a Bernoulli reformulation of the DC-SBM algorithm as the basis for our generator (Rohe et al., 2018). Its justification and precise correspondence with the Poisson DC-SBM are deferred to Appendix  B.
@@ 4.1 Three-Level Architecture
  • Structural patterns: The universe-level edge propensity matrix $\tilde{\mathbf{P}}\in\mathbb{R}^{K\times K}$ encodes relative inter-community connection strengths. Unlike standard DC-SBMs with uniform block probabilities, we introduce heterogeneity by generating $\tilde{P}_{rs}=1+\xi_{rs}$ where $\xi_{rs}\sim\mathcal{N}(0,(2\epsilon)^{2})$ with variance parameter $\epsilon$. This perturbation is symmetrized and rescaled to preserve target homophily and degree constraints (details in App. C), yielding fine-grained structural variation across community pairs.
@@
4.1.1 Graph Instance Generation Process
@@
The resulting matrix $\mathbf{P}_{\mathrm{scaled}}$ preserves the heterogeneity of $\mathbf{P}_{\mathrm{sub}}$ while satisfying both constraints (see Appendix C).
@@
    Degree factors. For each node $i$ in community $c(i)$, we assign a degree factor $\theta_{i}$ by coupling degree distributions to communities. We first sample power-law degree factors and sort them as $(\theta_{(1)},\dots,\theta_{(n)})$. Each community’s degree center $\delta_{c(i)}$ maps to a preferred rank $\mu_{c(i)}=\frac{1+\delta_{c(i)}}{2}(n-1)$, and we assign degree factors by sampling rank indices from $\mathcal{N}(\mu_{c(i)},\sigma^{2})$ truncated to $[1,n]$, where $\sigma^{2}$ is determined by the degree separation parameter $\rho$. Full details are in Appendix D.
@@
    After sampling, we verify connectivity and connect any disconnected components by adding edges that minimize deviation from the target block structure $\mathbf{P}_{\mathrm{scaled}}$ (details in Appendix E).
@@ 4.2 Validation of GraphUniverse
To ensure our multiple graph generation framework produces high-fidelity graphs with the intended properties and learnable signals within and across graphs, we conduct a comprehensive validation study examining how generation parameters affect three critical aspects: graph properties, signal strength, and cross-graph consistency. This systematic analysis serves both to validate the correctness of our implementation and to characterize the parameter space for downstream applications.
@@
We define three categories of validation metrics. Graph Property metrics verify that generated graphs match target structural characteristics, including homophily levels, average degree, degree distribution tails, and deviations from expected community edge probability matrices. Signal Strength metrics assess the predictability of community labels using different node-level features (node features, degree, and multi-hop neighborhood structure), ensuring graphs contain learnable signals for downstream tasks. Cross-Graph Consistency metrics evaluate whether community identities remain semantically consistent across different graph instances through feature centroid similarity, structural pattern correlation, and degree ranking preservation. Detailed definitions and implementation details for all metrics are provided in Appendix G.
@@
We generate 100 distinct graph families—30 graphs each—with completely randomized parameter configurations sampled uniformly across broad ranges (further details in Appendix F). This stress-tests the framework’s robustness by capturing parameter interactions at extreme values rather than nominal operating points. In practice, to do so we assess parameter-metric relationships using Pearson correlations on family-level means, reporting only statistically significant correlations ($p<0.05$). The correlation heatmap of Figure 2 shows the parameter (y-axis) responsiveness across all validation metrics (x-axis).
@@
Graph property metrics (left panel in 2) show expected strong correlations: parameters precisely control homophily and average degree, with slight deviations under extreme settings reflecting our multiplicative edge generation process. Signal strength metrics (middle panel in 2) confirm theoretical expectations: cluster variance controls feature signals, homophily/degree enhance structure signals, and fewer communities simplify classification. Cross-graph consistency metrics (right panel in 2) provide the strongest validation of our hierarchical design, with propensity variance governing structure consistency, degree separation controlling degree consistency, and cluster variance determining feature consistency. A detailed analysis of the parameter validation results, as well as individual parameter effect plots, are provided in Appendix Section H.
This comprehensive analysis demonstrates that our generation framework successfully translates user-specified parameters into measurable, controllable graph properties without significant unexpected correlations, validating its suitability for systematic graph learning evaluation.
@@ 4.3 Scalability
Avg. number Time per Throughput R-squared of nodes graph (sec) (graphs/sec) (graphs) 100 0.020 45.5 0.999 1000 1.070 0.9 0.998
GraphUniverse demonstrates linear scaling across graph sizes, enabling efficient large-scale evaluation. Table 1 shows generation performance measured on an AMD Ryzen 7 5700U CPU processor (single-threaded), averaged over 500 graphs.
@@ 4.4 Interactive Exploration Tool
To complement the open-source package, we also provide an interactive Streamlit application for code-free exploration of the GraphUniverse generator (https://graphuniverse.streamlit.app/). The app allows users to define a generation universe, tune the family parameters (e.g., homophily, degree distribution), and instantly visualize the resulting graphs and their properties—as well as the validation metrics confirming that the dataset conforms to the desired properties. Furthermore, generated graph families can be downloaded as a PyTorch Geometric InMemoryDataset object, enabling a seamless transition from interactive design to local model training and experimentation.
@@ 5.1 Experimental Setup
All experiments are conducted using the TopoBench benchmarking framework (Telyatnikov et al., 2024), which we extend with a custom GraphUniverse loader to systematically define and iterate over graph generation parameters. [^1] Code and config. files to reproduce all experiments will be made publicly available upon acceptance. Across our benchmarking, we evaluate models in both inductive and transductive settings on two distinct tasks: node-level community detection (classification) and graph-level triangle counting (regression). Unless otherwise specified, we consider a set of fixed dataset generation parameters—such as the number of graphs in inductive settings (1000), or the number of nodes in transductive ones (1000)—to ensure a consistent baseline across experiments (defaults lists in Appendix I.1).
@@
We evaluate a diverse set of architectures representing major paradigms in contemporary graph learning (a brief description of each of them can be found in Appendix I.2):
  • Non message-passing baselines: DeepSet (Zaheer et al., 2017) and GraphMLP (Hu et al., 2021).
  • Message-passing GNNs: GCN (Kipf, 2016), GraphSAGE (Hamilton et al., 2017), GIN (Xu et al., 2018) and GATv2 (Brody et al., 2021).
  • Topological Models: TopoTune (Papillon et al., 2025) and Neural Sheaf Diffusion (Bodnar et al., 2022).
  • Graph Transformers: GPS (GraphGPS with positional encodings) (Rampášek et al., 2022).
@@
For each model-dataset configuration, we conduct comprehensive grid search hyperparameter optimization using architecture-specific parameter grids detailed in the Appendix I.3. Each configuration is evaluated across 3 different dataset instantiations generated with the same input parameters but different data random seeds, with the configuration achieving the highest mean validation performance selected for final test evaluation.
@@
We systematically vary three fundamental graph properties across families of 1000 graphs each: homophily range ($[0.0,0.1]$, $[0.4,0.6]$, $[0.9,1.0]$), average degree range ($[1,5]$, $[5,10]$, $[10,20]$), and cluster variance (basically feature noise, setting as options $0.2$, $0.5$ and $0.8$), keeping all other graph family parameters at default values. We directly compare these inductive results against equivalent single-graph transductive evaluation using GraphWorld-style generation (with identical property values to quantify the performance gap between transductive and inductive settings). Figure 3.A reports on both the model’s test accuracy.
Figure 3: A) Inductive (graph families of 1000 graphs) versus transductive (single graphs) test accuracy on community detection across different graph properties, with each architecture individually optimized. B) Distribution shift analysis: best-performing inductive models evaluated on graph families with shifted properties from the same Universe. Plots show accuracy changes under distributional shifts, with x-axis indicating the original training domain.
@@
Insight 3: Performance Directly Correlates with Feature Quality.
As expected, increasing cluster variance—a parameter that only injects noise into node features without affecting graph structure—consistently degrades performance for all models.

### v2

# arXiv:2509.21097v2[cs.LG] 28 Feb 2026

@@ Abstract
A fundamental challenge in graph learning is understanding how models generalize to new, unseen graphs. While synthetic benchmarks offer controlled settings for analysis, existing approaches are confined to single-graph, transductive settings where models train and test on the same graph structure. Addressing this gap, we introduce GraphUniverse, a framework for generating entire families of graphs to enable the first systematic evaluation of inductive generalization at scale. Our core innovation is the generation of graphs with persistent semantic communities, ensuring conceptual consistency while allowing fine-grained control over structural properties like homophily and degree distributions. This enables crucial but underexplored robustness tests, such as performance under controlled distribution shifts. Benchmarking a wide range of architectures—from GNNs to graph transformers and topological architectures—reveals that strong transductive performance is a poor predictor of inductive generalization. Furthermore, we find that robustness to distribution shift is highly sensitive not only to model architecture choice but also to the initial graph regime (e.g., high vs. low homophily). Beyond benchmarking, GraphUniverse’s flexibility and scalability can facilitate the development of robust and truly generalizable architectures. The framework is open-source at https://github.com/LouisVanLangendonck/GraphUniverse.
@@ 1 Introduction
Graph learning has emerged as a powerful paradigm for learning from relational data across diverse domains, from drug discovery (Wong et al., 2024) and fraud detection (Cheng et al., 2025) to knowledge graphs (Galkin et al., 2023). Graph Neural Networks (GNNs) (Scarselli et al., 2008), with its countless variants (Gilmer et al., 2017; Kipf and Welling, 2017; Hamilton et al., 2017; Xu et al., 2019; Veličković et al., 2018), have demonstrated remarkable success in extending deep learning frameworks to graph-structured data, achieving competitive performance on tasks ranging from node classification to graph-level prediction. However, unlike the transformative leap from task-specific models to general-purpose architectures observed in natural language processing and computer vision, graph learning remains largely limited to specialized, task-specific models with limited evidence of robust generalization and scaling capabilities.
@@
    We develop a hierarchical generative model extending Degree Corrected-Stochastic Block Models (DC-SBMs) (Karrer and Newman, 2011) to an inductive setting with multiple graphs that maintain semantic consistency—i.e. node identities or community structures persist across different graph instances—while enabling controlled variation in their structural properties.
@@
    GraphUniverse is available as a PyPI package at https://pypi.org/project/graph-universe/, with source code at https://github.com/LouisVanLangendonck/GraphUniverse. We have integrated GraphUniverse into TopoBench (Telyatnikov et al., 2025) to facilitate reproduction and extension of our experimental results.
  3. 3.
    We conduct systematic benchmarking comparing inductive and transductive evaluation across classical and contemporary graph architectures, revealing differences in model rankings between paradigms. Additionally, we evaluate model robustness under controlled property shifts, an analysis only possible with our inductive framework, finding that robustness strongly depends on both architecture choice and initial graph properties. These findings challenge conventional assumptions about graph model performance and demonstrate the critical importance of evaluation paradigm choice in assessing true model capabilities. Furthermore, we demonstrate that GraphUniverse-generated datasets can serve as effective proxies for real-world datasets, with model rankings showing strong correlations with those obtained on real data.
  4. 4.
@@
We envision GraphUniverse as a versatile tool for diverse research applications, from targeted generalization benchmarks to large-scale data generation and augmentation for model pre-training. While our experiments demonstrate its immediate utility, they represent only a fraction of what is possible with controllable graph family generation. Thus, we release GraphUniverse to the community as a flexible framework, inviting extensions and adaptations to explore new frontiers in graph learning.
@@ 2 Related Work
Recognizing the limitations of static datasets, a growing consensus advocates for high-fidelity synthetic data generation as a path toward more principled and scalable evaluation (Bechler-Speicher et al., 2025; Wang et al., 2025). The most prominent effort in this direction is GraphWorld (Palowitch et al., 2022), which uses synthetic generation to study model performance across a space of graph properties. Related efforts like CGT (Zahirnia et al., 2023) and the metadata-driven approach (Li et al., 2023) provide valuable insights into model behavior by mapping real-world graphs to synthetic equivalents and analyzing performance across graph properties, respectively. While these approaches enable controlled analysis, they remain confined to generating independent, single graphs. This restricts evaluation to the transductive setting, where models are tested on the same graph structure seen during training, and fundamentally prevents the study of a model’s ability to generalize to entirely new and unseen graphs.
The importance of synthetic data extends beyond benchmarking to foundation model development. GraphFM (Lachi et al., 2024) leverages GraphWorld-style synthetic graphs to expand its pretraining corpus, though limited to transductive settings, while OpenGraph (Xia et al., 2024) uses LLMs to augment existing datasets—mirroring synthetic data’s established role in computer vision (Tobin et al., 2017) and NLP (Wang et al., 2023). However, existing graph generation frameworks cannot provide the inductive generalization and systematic coverage of graph modalities that robust foundation models require (Bechler-Speicher et al., 2025; Wang et al., 2025)—a capability limited to multi-graph approaches.
GraphUniverse directly addresses these limitations. Unlike static benchmarks, it provides a generative framework capable of producing unlimited data with fine-grained control over structural properties. Critically, unlike GraphWorld and other single-graph approaches, it generates entire families of graphs with shared semantic meaning, enabling systematic study of inductive generalization. Furthermore, our experiments validate that GraphUniverse-generated data closely mirrors real-world model behavior, suggesting its potential as a complementary data source for model development, including pre-training applications. We present GraphUniverse as an open-source tool that researchers can extend and adapt for their specific needs, whether for controlled benchmarking or as a basis for more sophisticated data augmentation strategies for Graph Foundation Model developement, of which a detailed discussion is provided in Appendix B.
@@ 3 Background
A limitation of SBM is that it enforces homogeneous expected degrees within a community, since edge probabilities depend only on block membership. To address this, the original DC-SBM (Karrer and Newman, 2011) implementation introduces node-specific degree factors $\theta_{i}>0$ and a nonnegative block matrix $\Lambda\in\mathbb{R}_{\geq 0}^{k\times k}$. Moreover, they shift focus from simple graphs to multigraphs, i.e. a graph with multi-edges and self-loops. In its original Poisson multigraph form, edges are counts with rates:
@@
so $\Lambda$ controls the total number of edges in a block, while $\theta$ redistributes degrees within blocks. This Poisson formulation is standard in inference-focused works (Karrer and Newman, 2011; Abbe, 2018).
@@
Since the original DC-SBM naturally generate Poisson multigraphs, a common approach to get simple graphs from them is to collapse multi-edges into single undirected edges and remove self-edges after generation (as done in GraphWorld (Palowitch et al., 2022)). However, this leads to a systematic but unpredictable mismatch between the input parameters and the properties of the output graph. Therefore, we choose to work directly with a Bernoulli reformulation of the DC-SBM algorithm as the basis for our generator (Rohe et al., 2018). Its justification and precise correspondence with the Poisson DC-SBM are deferred to Appendix  C. Moreover, a discussion of the limitations of relying on DC-SBM models (e.g. lack of higher order structures) is provided in Appendix D.
@@ 4.1 Three-Level Architecture
  • Structural patterns: The universe-level edge propensity matrix $\tilde{\mathbf{P}}\in\mathbb{R}^{K\times K}$ encodes relative inter-community connection strengths. Unlike standard DC-SBMs with uniform block probabilities, we introduce heterogeneity by generating $\tilde{P}_{rs}=1+\xi_{rs}$ where $\xi_{rs}\sim\mathcal{N}(0,(2\epsilon)^{2})$ with variance parameter $\epsilon$. This perturbation is symmetrized and rescaled to preserve target homophily and degree constraints (details in App. E), yielding fine-grained structural variation across community pairs.
@@
4.2 Graph Instance Generation Process
@@
The resulting matrix $\mathbf{P}_{\mathrm{scaled}}$ preserves the heterogeneity of $\mathbf{P}_{\mathrm{sub}}$ while satisfying both constraints (see Appendix E).
@@
    Degree factors. For each node $i$ in community $c(i)$, we assign a degree factor $\theta_{i}$ by coupling degree distributions to communities. We first sample power-law degree factors and sort them as $(\theta_{(1)},\dots,\theta_{(n)})$. Each community’s degree center $\delta_{c(i)}$ maps to a preferred rank $\mu_{c(i)}=\frac{1+\delta_{c(i)}}{2}(n-1)$, and we assign degree factors by sampling rank indices from $\mathcal{N}(\mu_{c(i)},\sigma^{2})$ truncated to $[1,n]$, where $\sigma^{2}$ is determined by the degree separation parameter $\rho$. Full details are in Appendix F.
@@
    After sampling, we verify connectivity and connect any disconnected components by adding edges that minimize deviation from the target block structure $\mathbf{P}_{\mathrm{scaled}}$ (details in Appendix G).
@@ 4.3 Validation of GraphUniverse
To ensure our multiple graph generation framework produces high-fidelity graphs with the intended properties and learnable signals within and across graphs, we conduct a comprehensive validation study into three critical aspects: graph properties, signal strength, and cross-graph consistency.
@@
We define three categories of validation metrics. Graph Property metrics verify that generated graphs match target structural characteristics, including homophily levels, average degree, degree distribution tails, and deviations from expected community edge probability matrices. Signal Strength metrics assess the predictability of community labels using different node-level features (node features, degree, and multi-hop neighborhood structure), ensuring graphs contain learnable signals for downstream tasks. Cross-Graph Consistency metrics evaluate whether community identities remain semantically consistent across different graph instances through feature centroid similarity, structural pattern correlation, and degree ranking preservation. Detailed definitions for all metrics are provided in Appendix I.
@@
We generate 100 distinct graph families—30 graphs each—with completely randomized parameter configurations sampled uniformly across broad ranges (further details in Appendix H). This stress-tests the framework’s robustness by capturing parameter interactions at extreme values rather than nominal operating points. We assess parameter-metric relationships using Pearson correlations on family-level means, reporting only statistically significant correlations ($p<0.05$). The correlation heatmap of Figure 2 shows the parameter (y-axis) responsiveness across all validation metrics (x-axis).
@@
Graph property metrics (left panel in 2) show expected strong correlations: parameters precisely control homophily and average degree, with slight deviations under extreme settings reflecting our multiplicative edge generation process. Signal strength metrics (middle panel in 2) confirm theoretical expectations: cluster variance controls feature signals, homophily/degree enhance structure signals, and fewer communities simplify classification. Cross-graph consistency metrics (right panel in 2) provide the strongest validation of our hierarchical design, with propensity variance governing structure consistency, degree separation controlling degree consistency, and cluster variance determining feature consistency. A detailed analysis of the parameter validation results, as well as individual parameter effect plots, are provided in Appendix Section J.
@@ 4.4 Scalability
Avg. number Time per Throughput of nodes graph (sec) (graphs/sec) 10 0.002 449.7 100 0.023 42.8 500 0.349 2.9 1000 1.309 0.8
GraphUniverse demonstrates linear scaling across graph sizes, enabling efficient large-scale evaluation. Table 1 shows generation performance measured on an AMD Ryzen 7 5700U CPU processor (single-threaded), averaged over 100 graphs.
@@ 4.5 Interactive Exploration Tool
To complement the open-source package, we provide an interactive Streamlit application for code-free exploration of the GraphUniverse generator, available both as a hosted demo (https://graphuniverse.streamlit.app/) and for local deployment via the PyPI package. The app allows users to define a generation universe, tune family parameters (e.g., homophily, degree distribution), and instantly visualize the resulting graphs, their properties and validation metrics. Generated graph families can be downloaded as PyTorch Geometric InMemoryDataset objects.
@@ 5.1 Experimental Setup
All experiments are conducted using the TopoBench benchmarking framework (Telyatnikov et al., 2025), which we extend with a custom GraphUniverse loader to systematically define and iterate over graph generation parameters. Across our benchmarking, we evaluate models in both inductive and transductive settings on two distinct tasks: node-level community detection (classification) and graph-level triangle counting (regression). Unless otherwise specified, we consider a set of fixed dataset generation parameters—such as the number of graphs in inductive settings (1000), or the number of nodes in transductive ones (1000)—to ensure a consistent baseline across experiments (defaults lists in Appendix M.1).
@@
We evaluate a diverse set of architectures representing major paradigms in contemporary graph learning (a brief description of each of them can be found in Appendix M.2): DeepSet (Zaheer et al., 2017), GraphMLP (Hu et al., 2021), GCN (Kipf and Welling, 2017), GraphSAGE (Hamilton et al., 2017), GIN (Xu et al., 2019), GATv2 (Brody et al., 2022), TopoTune (Papillon et al., 2025), Neural Sheaf Diffusion (Bodnar et al., 2022), and GPS (Rampášek et al., 2022) (see Appendix M.2).
@@
For each model-dataset configuration, we conduct comprehensive grid search hyperparameter optimization using architecture-specific parameter grids detailed in the Appendix M.3. Each configuration is evaluated across 3 different dataset instantiations generated with the same input parameters but different data random seeds, with the configuration achieving the highest mean validation performance selected for final test evaluation.
@@
We systematically vary three fundamental graph properties across families of 1000 graphs each: homophily range ($[0.0,0.1]$, $[0.4,0.6]$, $[0.9,1.0]$), average degree range ($[1,5]$, $[5,10]$, $[10,20]$), and cluster variance (basically feature noise, setting as options $0.2$, $0.5$ and $0.8$), keeping all other graph family parameters at default values. We directly compare these inductive results against equivalent single-graph transductive evaluation using GraphWorld-style generation (with mean property of corresponding inductive setting). Figure 3.A reports on the results. Appendix L extends the homophily analysis to specific heterophilic GNN architectures.
Figure 3: A) Inductive (graph families of 1000 graphs) versus transductive (single graphs) test accuracy on community detection across different graph properties, with each architecture individually optimized. B) Distribution shift analysis: best-performing inductive models evaluated on graph families with shifted properties from the same Universe. Plots show accuracy changes under distributional shifts, with x-axis indicating the original training domain. N/A indicates shifts beyond feasible parameter bounds.